OpenStudy (anonymous):

I need help determining the fractional percentage of a number. Ie. 4% of 36 = ? Please show your steps.

jigglypuff314 (jigglypuff314):

Hello and Welcome to Openstudy! :) 4% in fraction form is 4 over 100 and "of" means multiplying so multiply 4/100 * 36 =

OpenStudy (anonymous):

there is a formula : is/of = %/100 now read your problem, you will be given 2 numbers and you just have to find the third. 4% of 36 = (= is the same as is) is/of = %/100 your are given the percent , it is 4 you are given the " of " term, it is 36 so we are looking for the " is " term which we will use x to represent that term now we sub into our formula.. x/36 = 4/100 (this is a proportion and we need to cross multiply) (100)(x) = (36)(4) 100x = 144 -- divide by 100 x = 144/100 = 1.44 4% of 36 = 1.44

OpenStudy (anonymous):

Made a mistake typing the question... it is .4% or 0.4%.

OpenStudy (calculusxy):

Remember that a percent is over a 100. A decimal is usually written upon a hundred. And writing the decimal (if it is in the hundredths place), then the fraction will be over a Hundred. Practically, we use 100 as our base. Are you with me so tar?

OpenStudy (calculusxy):

*far

OpenStudy (anonymous):

yea

OpenStudy (calculusxy):

Since we have .4 % then that means that we write that as a decimal. Since to convert a decimal to percent, we move the decimal place two places to the RIGHT. Now since we are doing the inverse, you need to go to the LEFT. If you move the decimal of 0.4 two places the left, we get 0.004. Now we want to find what is 0.004 (decimal) of 36. That means we multiply 0.004 x 36 to get our answer. Also remember that 'of' means to multiply in math when it is between two numbers.

OpenStudy (calculusxy):

Do you get how to do this? If you have any confusions, tell me so that I can help you get rid of it.

OpenStudy (anonymous):

That makes sense, Thanks :)

OpenStudy (calculusxy):

If you like my answer, you can be my fan or give me a medal :)

OpenStudy (calculusxy):

Have a good day :)